Problem 1-21 Traditional and Contribution Format Income Statements [LO1-6]

Marwick’s Pianos, Inc., purchases pianos from a large manufacturer for an average cost of $1,480 per unit and then sells them to retail customers for an average price of $2,900 each. The company’s selling and administrative costs for a typical month are presented below:

Costs Cost Formula
Selling:
Advertising $ 948 per month
Sales salaries and commissions $ 4,830 per month, plus 5% of sales
Delivery of pianos to customers $ 61 per piano sold
Utilities $ 652 per month
Depreciation of sales facilities $ 5,075 per month
Administrative:
Executive salaries $ 13,584 per month
Insurance $ 709 per month
Clerical $ 2,545 per month, plus $39 per piano sold
Depreciation of office equipment $ 917 per month

During August, Marwick’s Pianos, Inc., sold and delivered 61 pianos.

Required:

1. Prepare a traditional format income statement for August.
2. Prepare a contribution format income statement for August. Show costs and revenues on both a total and a per unit basis down through contribution margin.
Answer #1

req1 B F HIJ K req 1 req 2 $ 1,76,900 $ 90,280 $ 86,620 Marwicks Pianos, Inc. Traditional Income Statement For the Month of

For formulas and calculations, refer to the image below -

X fic req1 B E F req 1 reg 2 =2900*61 =1480*61 =E5-E6 61 Marwicks Pianos, Inc. Traditional Income Statement For the Month of

In case you have any query, kindly ask in comments.